# Cost-Aware Contrastive Routing for LLMs

### v1

@@ A.3 Routing as Recommendation
Routing can also be framed as a recommendation problem, wherein the input query plays the role of a "user," the pool of LLMs corresponds to "items," and past performance metrics form the implicit interaction history Zhao et al. (2024a); Wu et al. (2022). However, unlike conventional recommender systems, routing has limited "user" features (i.e., input metadata), making label collection and generalization especially challenging Nguyen et al. (2024); Liu et al. (2024).
Matrix factorization, attention-based models, and graph neural networks are used in both recommenders and routers Ong et al. (2024); Zhuang et al. (2024); Feng et al. (2024), reinforcing the close link between the two domains.
@@ A.4 Scaling Laws and Architecture Trends
Scaling laws Kaplan et al. (2020); Li et al. (2025) describe predictable trends between model size, data, and performance, guiding the development of efficient LLM architectures. These insights have been extended to MoEs Fedus et al. (2022); Clark et al. (2022), sparse models Frantar et al. (2023), and hybrid systems Gu and Dao (2023); Poli et al. (2024), offering context for when routing or merging approaches might be most beneficial.
@@ A.5 Routing within MoE and Hybrid Architectures
Routing is a central mechanism in MoE models Jacobs et al. (1991); Jordan and Jacobs (1993); Shazeer et al. (2017), where expert modules are dynamically activated based on input. While classical MoEs involved equally-sized sub-models, modern approaches like Switch Transformer Fedus et al. (2022) and Mixtral Jiang et al. (2024) employ sparse activation to minimize cost.
Routing LLMs can be viewed as a coarse-grained MoE, where each expert is a full LLM. Approaches like UltraFuser Ding et al. (2024b), Branch-Train-MiX Sukhbaatar et al. (2024), and token-level fusion highlight recent advances in combining model specialization and flexibility.
@@ B.1 Notation and Preliminaries
Let the Bayes-optimal router be $r^{\star}(x)=\mathop{\rm arg\,min}_{m\in[M]}\bigl{[}\gamma(x,h_{m})+\lambda c_{m}\bigr{]}$ and define the excess cost–error gap
@@
Hence $\Delta_{r^{\star}(x)}(x)=0$ and $\Delta_{m}(x)\geq 0$. For a query $x$ let $r_{k}(x)$ be the radius of the ball $B(x,r)\subset\mathbb{S}^{d-1}$ (in the cosine metric) that contains the $k$-th nearest training neighbor. If the marginal on $q(\mathcal{X})$ has a density, $\operatorname*{\mathbb{E}}\!\bigl{[}r_{k}(x)^{d}\bigr{]}\leq C_{1}k/n$ Cover and Hart (1967); Stone (1977).
@@
Let the empirical cost-adjusted risk be
@@
For convenience write $S_{im}=q_{i}^{\top}e_{m}/\tau_{k}$ when $m\in\mathcal{B}_{k}$. The single-query loss (Equation (8)) is
@@
Taking the gradient w.r.t. $q_{i}$ and setting it to zero gives $\sum_{m}(p_{im}-q_{im})e_{m}=0$. Project onto $q_{i}$: $\sum_{m}(p_{im}-q_{im})S_{im}=0.$ Fix $m^{+}\in\mathcal{P}_{ik}$, $m^{-}\in\mathcal{N}(i)$ with $c_{m^{+}}\leq c_{m^{-}}$. Because $p_{im^{-}}=0$ while $p_{im^{+}}>0$, the equality forces $q_{im^{+}}>q_{im^{-}}$, hence $S_{im^{+}}-\gamma c_{m^{+}}>S_{im^{-}}-\gamma c_{m^{-}}$. Re-arranging yields $q_{i}^{\top}e_{m^{+}}>q_{i}^{\top}e_{m^{-}}$, establishing directional alignment.
@@ C.1 Logit-Footprint Descriptors
These tokens aren’t used for their meaning. They’re probes of each model’s output behavior. Even common words get scored differently across models (temperature, punctuation/number handling, style). By averaging over many prompts and steps, the descriptor captures overall model behavior, not any single word’s semantics. Also, a recent work Sun et al. (2025) shows that LLMs exhibit stable, word-level idiosyncrasies (as the authors call them) that enable near-perfect model attribution using only the first few generated tokens (even after paraphrasing or translation), implying that common tokens still provide discriminative signals about a model’s predictive calibration.
@@ C.4 Dense Human Annotations
  • Positive sets $P(i)$ with preference structure. Replace binary "correct expert" labels with pairwise preferences (cheap-and-good $\succ$ expensive-and-similar $\succ$ clearly wrong), yielding band-aware positives and margin constraints. This can be implemented by expanding $P(i)$ and adding a lightweight pairwise ranking (DPO-style) regularizer within each cost band.
@@ D.1 Experimental Settings
We compare our proposed method against a comprehensive set of baselines designed to capture key routing strategies and their trade-offs. Specifically, we include UMR Jitkrittum et al. (2025), a recent state-of-the-art technique that clusters prompt embeddings to route queries to LLM pools efficiently; Thompson Sampling Li (2025); Agrawal and Goyal (2012), which frames routing as a bandit exploration-exploitation problem to balance cost and accuracy dynamically; Pareto-optimal routing Hu et al. (2024), a strategy that selects models by explicitly considering the cost-accuracy Pareto frontier; and two extreme baselines—Random, which selects models uniformly at random to represent naive routing without intelligent selection, and Oracle (Clairvoyant Upper-Bound Jitkrittum et al. (2025)), which always selects the most accurate model at the lowest possible cost and thus represents a theoretical performance ceiling. Additionally, we evaluate against parametric gating methods (Parametric Softmax Router) inspired by classical mixture-of-experts architectures Ong et al. (2024) and SoftMoE, which models router decisions via differentiable soft gating functions Puigcerver et al. (2024). Collectively, these baselines enable us to rigorously assess whether our contrastive routing approach delivers meaningful improvements in performance, cost-efficiency, and generalization capabilities relative to existing strategies.
@@
We train our router and evaluate it on three datasets: EmbedLLM Zhuang et al. (2024), MixInstruct Jiang et al. (2023), and RouterBench Hu et al. (2024). For EmbedLLM and MixInstruct, we sample 192 probes from their respective validation sets. Each probe is processed to extract logit-based descriptors by capturing the top $K=256$ tokens over a horizon of $T=10$ tokens (Equation (2)), resulting in a 256-dimensional vector per model. For RouterBench, we sample 192 probes from its training set, ensuring these probes are excluded from the training data used for the contrastive router. We compute perplexity-based descriptors on RouterBench and use GPT-2 Radford et al. (2019). On both EmbedLLM and RouterBench, we use binary accuracy as the per-sample evaluation metric, meaning an LLM response is classified strictly as correct or incorrect. For MixInstruct, we employ exponentiated BARTScore Yuan et al. (2021) as the evaluation metric, following the approach in prior work Jitkrittum et al. (2025); Jiang et al. (2023).
We evaluate each routing strategy using a deferral curve Jitkrittum et al. (2025) which plots the average response quality against the total inference cost. Sweeping the routing penalty parameter $\lambda$ over the interval $\lambda\!\in\![0,\lambda_{\max}]$ (refer to Equation (9)) traces the deferral curve. For the EmbedLLM and MixInstruct datasets, we define the cost of processing a prompt as the number of parameters in the LLM, serving as a proxy for computational resources and latency. In the case of RouterBench, we utilize the actual API call costs in USD, as provided in the dataset. Following Jitkrittum et al. (2025) we employ evaluation metrics including Area Under the Deferral Curve (AUDC), Query-Normalized Cost (QNC), and peak accuracy. QNC is the minimum relative cost required to match the performance of the most accurate tested LLM.
@@
We use a frozen sentence-transformers/all-MiniLM-L6-v2Reimers and Gurevych (2019) model as the embedding backbone ($\Phi(x)$ in Section3) across all experiments. Our trainable router component is a two-layer MLP, denoted as $g_{\theta}(.)$, which projects prompt embeddings into the expert descriptor space. We train our contrastive router on the training splits of each dataset, excluding the probe examples from RouterBench. Training is performed for 10 epochs using the AdamW optimizer with a batch size of 512 and a learning rate of $5\times 10^{-4}$. For the cost spectrum loss (Equation (8)), we set the number of cost bands to $K=5$ and the negative cost penalty to $\lambda=0.1$. The hyperparameters for the linear schedule of band-specific temperatures (Equation (7)) are set as $\alpha=0.25$ and $\tau_{\min}=0.05$. All training and descriptor extraction are done on RTX6000Ada GPUs with $\sim$48GB GPU memory.
@@ D.3.1 Ablation of Number of Cost bands
Figure 7: Deferral-curve metrics across different numbers of cost bands.
@@
Table 8: Deferral-curve metrics across different numbers of cost bands.
@@ D.3.2 Ablation of the Number of Neighbors
Increasing $k$ (the number of ANN neighbors selected before cost-aware scoring) generally provides modest improvements before reaching a plateau. (The baseline UMR Jitkrittum et al. (2025) also included an ablation on $k$ in a K-NN router.) Beyond a small $K$, the gains in accuracy or AUDC become minimal, while both latency and the likelihood of choosing unnecessarily expensive experts increase. We selected a default of $K=4$ and it worked reasonably well so we did not do further hyperparameter tuning. We performed an ablation on a subset of embedllm prompts. The results are shown in Table 9 which are consistent with the trend observed in UMR.
@@ D.6 Statistical Significance
We perform full evaluation on Embedllm Zhuang et al. (2024) using paired, prompt-level significance tests to concretely assess statistical significance. Specifically, we computed a paired bootstrap Efron and Tibshirani (1994) (sampling 3,000 prompts with replacement, 5,000 times) to obtain a 95% confidence interval for $\Delta_{\mathrm{AUDC}}=\mathrm{AUDC}_{\text{CSCR}}-\mathrm{AUDC}_{\text{UMR}}$ (UMR is the best baseline overall). We also report a one-sided p-value for the hypothesis $\Delta>0$. Additionally, we ran McNemar’s McNemar (1947) test at a matched budget (using the median of the combined cost grids) to compare per-prompt wins and losses at equivalent operating cost. These tests quantify uncertainty over the test prompts.

### v2

@@ A.3 Routing as Recommendation
Routing can also be framed as a recommendation problem, wherein the input query plays the role of a "user," the pool of LLMs corresponds to "items," and past performance metrics form the implicit interaction history [100, 93]. However, unlike conventional recommender systems, routing has limited "user" features (i.e., input metadata), making label collection and generalization especially challenging [61, 49].
Matrix factorization, attention-based models, and graph neural networks are used in both recommenders and routers [62, 105, 25], reinforcing the close link between the two domains.
@@ A.4 Scaling Laws and Architecture Trends
Scaling laws [42, 47] describe predictable trends between model size, data, and performance, guiding the development of efficient LLM architectures. These insights have been extended to MoEs [24, 13], sparse models [26], and hybrid systems [28, 63], offering context for when routing or merging approaches might be most beneficial.
@@ A.5 Routing within MoE and Hybrid Architectures
Routing is a central mechanism in MoE models [36, 41, 75], where expert modules are dynamically activated based on input. While classical MoEs involved equally-sized sub-models, modern approaches like Switch Transformer [24] and Mixtral [38] employ sparse activation to minimize cost.
Routing LLMs can be viewed as a coarse-grained MoE, where each expert is a full LLM. Approaches like UltraFuser [20], Branch-Train-MiX [81], and token-level fusion highlight recent advances in combining model specialization and flexibility.
@@ B.1 Notation and Preliminaries
Let the Bayes‑optimal router be $r^{\star}(x)=\mathop{\rm arg\,min}_{m\in[M]}\bigl[\gamma(x,h_{m})+\lambda c_{m}\bigr]$ and define the excess cost–error gap
@@
Hence $\Delta_{r^{\star}(x)}(x)=0$ and $\Delta_{m}(x)\geq 0$. For a query $x$ let $r_{k}(x)$ be the radius of the ball $B(x,r)\subset\mathbb{S}^{d-1}$ (in the cosine metric) that contains the $k$-th nearest training neighbor. If the marginal on $q(\mathcal{X})$ has a density, $\operatorname*{\mathbb{E}}\!\bigl[r_{k}(x)^{d}\bigr]\leq C_{1}k/n$ [15, 79].
@@
Let the empirical cost‑adjusted risk be
@@
For convenience write $S_{im}=q_{i}^{\top}e_{m}/\tau_{k}$ when $m\in\mathcal{B}_{k}$. The single‑query loss (Equation (8)) is
@@
Taking the gradient w.r.t. $q_{i}$ and setting it to zero gives $\sum_{m}(p_{im}-q_{im})e_{m}=0$. Project onto $q_{i}$: $\sum_{m}(p_{im}-q_{im})S_{im}=0.$ Fix $m^{+}\in\mathcal{P}_{ik}$, $m^{-}\in\mathcal{N}(i)$ with $c_{m^{+}}\leq c_{m^{-}}$. Because $p_{im^{-}}=0$ while $p_{im^{+}}>0$, the equality forces $q_{im^{+}}>q_{im^{-}}$, hence $S_{im^{+}}-\gamma c_{m^{+}}>S_{im^{-}}-\gamma c_{m^{-}}$. Re‑arranging yields $q_{i}^{\top}e_{m^{+}}>q_{i}^{\top}e_{m^{-}}$, establishing directional alignment.
@@ C.1 Logit-Footprint Descriptors
These tokens aren’t used for their meaning. They’re probes of each model’s output behavior. Even common words get scored differently across models (temperature, punctuation/number handling, style). By averaging over many prompts and steps, the descriptor captures overall model behavior, not any single word’s semantics. Also, a recent work [82] shows that LLMs exhibit stable, word-level idiosyncrasies (as the authors call them) that enable near-perfect model attribution using only the first few generated tokens (even after paraphrasing or translation), implying that common tokens still provide discriminative signals about a model’s predictive calibration.
@@ C.4 Dense Human Annotations
  • Positive sets $P(i)$ with preference structure. Replace binary "correct expert" labels with pairwise preferences (cheap‐and‐good $\succ$ expensive‐and‐similar $\succ$ clearly wrong), yielding band-aware positives and margin constraints. This can be implemented by expanding $P(i)$ and adding a lightweight pairwise ranking (DPO-style) regularizer within each cost band.
@@ D.1 Experimental Settings
We compare our proposed method against a comprehensive set of baselines designed to capture key routing strategies and their trade-offs. Specifically, we include UMR [40], a recent state-of-the-art technique that clusters prompt embeddings to route queries to LLM pools efficiently; Thompson Sampling [48, 2], which frames routing as a bandit exploration-exploitation problem to balance cost and accuracy dynamically; Pareto-optimal routing [34], a strategy that selects models by explicitly considering the cost-accuracy Pareto frontier; and two extreme baselines—Random, which selects models uniformly at random to represent naive routing without intelligent selection, and Oracle (Clairvoyant Upper-Bound [40]), which always selects the most accurate model at the lowest possible cost and thus represents a theoretical performance ceiling. Additionally, we evaluate against parametric gating methods (Parametric Softmax Router) inspired by classical mixture-of-experts architectures [62] and SoftMoE, which models router decisions via differentiable soft gating functions [64]. Collectively, these baselines enable us to rigorously assess whether our contrastive routing approach delivers meaningful improvements in performance, cost-efficiency, and generalization capabilities relative to existing strategies.
@@
We train our router and evaluate it on three datasets: EmbedLLM [105], MixInstruct [39], and RouterBench [34]. For EmbedLLM and MixInstruct, we sample 192 probes from their respective validation sets. Each probe is processed to extract logit-based descriptors by capturing the top $K=256$ tokens over a horizon of $T=10$ tokens (Equation (2)), resulting in a 256-dimensional vector per model. For RouterBench, we sample 192 probes from its training set, ensuring these probes are excluded from the training data used for the contrastive router. We compute perplexity-based descriptors on RouterBench and use GPT-2 [66]. On both EmbedLLM and RouterBench, we use binary accuracy as the per-sample evaluation metric, meaning an LLM response is classified strictly as correct or incorrect. For MixInstruct, we employ exponentiated BARTScore [97] as the evaluation metric, following the approach in prior work [40, 39].
We evaluate each routing strategy using a deferral curve [40] which plots the average response quality against the total inference cost. Sweeping the routing penalty parameter $\lambda$ over the interval $\lambda\!\in\![0,\lambda_{\max}]$ (refer to Equation (9)) traces the deferral curve. For the EmbedLLM and MixInstruct datasets, we define the cost of processing a prompt as the number of parameters in the LLM, serving as a proxy for computational resources and latency. In the case of RouterBench, we utilize the actual API call costs in USD, as provided in the dataset. Following [40] we employ evaluation metrics including Area Under the Deferral Curve (AUDC), Query-Normalized Cost (QNC), and peak accuracy. QNC is the minimum relative cost required to match the performance of the most accurate tested LLM.
@@
We use a frozen sentence-transformers/all-MiniLM-L6-v2[70] model as the embedding backbone ($\Phi(x)$ in Section3) across all experiments. Our trainable router component is a two-layer MLP, denoted as $g_{\theta}(.)$, which projects prompt embeddings into the expert descriptor space. We train our contrastive router on the training splits of each dataset, excluding the probe examples from RouterBench. Training is performed for 10 epochs using the AdamW optimizer with a batch size of 512 and a learning rate of $5\times 10^{-4}$. For the cost spectrum loss (Equation (8)), we set the number of cost bands to $K=5$ and the negative cost penalty to $\lambda=0.1$. The hyperparameters for the linear schedule of band-specific temperatures (Equation (7)) are set as $\alpha=0.25$ and $\tau_{\min}=0.05$. All training and descriptor extraction are done on RTX6000Ada GPUs with $\sim$48GB GPU memory.
@@ D.3.1 Ablation of Number of Cost bands
Figure 7: Deferral‑curve metrics across different numbers of cost bands.
@@
Table 8: Deferral‑curve metrics across different numbers of cost bands.
@@ D.3.2 Ablation of the Number of Neighbors
Increasing $k$ (the number of ANN neighbors selected before cost-aware scoring) generally provides modest improvements before reaching a plateau. (The baseline UMR [40] also included an ablation on $k$ in a K-NN router.) Beyond a small $K$, the gains in accuracy or AUDC become minimal, while both latency and the likelihood of choosing unnecessarily expensive experts increase. We selected a default of $K=4$ and it worked reasonably well so we did not do further hyperparameter tuning. We performed an ablation on a subset of embedllm prompts. The results are shown in Table 9 which are consistent with the trend observed in UMR.
@@ D.6 Statistical Significance
We perform full evaluation on Embedllm [105] using paired, prompt-level significance tests to concretely assess statistical significance. Specifically, we computed a paired bootstrap [22] (sampling 3,000 prompts with replacement, 5,000 times) to obtain a 95% confidence interval for $\Delta_{\mathrm{AUDC}}=\mathrm{AUDC}_{\text{CSCR}}-\mathrm{AUDC}_{\text{UMR}}$ (UMR is the best baseline overall). We also report a one-sided p-value for the hypothesis $\Delta>0$. Additionally, we ran McNemar’s [57] test at a matched budget (using the median of the combined cost grids) to compare per-prompt wins and losses at equivalent operating cost. These tests quantify uncertainty over the test prompts.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We provide theoretical analysis in Section 5 and empirical results in Section 4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Provided in Section Limitations and Broader Impact
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide a theoretical analysis with proofs and assumptions in Section 5.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All experimental details are provided in Section 4 and  D.1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All datasets used are open source. We will provide the code for our experiments after paper decision is available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All experimental details are provided in Section 4 and  D.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We could not provide error bars deviation due to computational costs. No other prior work and baselines in our work do this either.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All experimental details are provided in Section 4 and  D.1. They include this information.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: There are not ethical concerns that we know of.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: A discussion of broader impact is provided in Section Limitations and Broader Impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All original owners of assets have been properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.